OpenStudy (anonymous):

x-intercept and y-intercept of the equation x + 2y = 4.

OpenStudy (anonymous):

For x-intercept: substitute y=0,x+2(0)=4,x=4 y-intercept:substitute x=0,0+2y=4,y=2

OpenStudy (anonymous):

x-intercept=4,y-intercept=2.

OpenStudy (anonymous):

thank you